OpenStudy (anonymous):

If x^3+ax+1=0 and x^4+ax^2+1=0 have a common root, then no. of values of a is ??

OpenStudy (anonymous):

assume that common root is \(t\) now subtract 2 equations and see what happens

OpenStudy (anonymous):

the equation u obtain from subtracting that 2 equation must has only one real root

OpenStudy (anonymous):

the eq obtained -

OpenStudy (anonymous):

try to factor it completely

OpenStudy (anonymous):

\[t^{4}-t^{3}+at^{2}-at=0\]??

OpenStudy (anonymous):

\( t^3(t-1)+at(t-1)=0 \)

OpenStudy (anonymous):

t^3+at=0

OpenStudy (anonymous):

t(t^2+a)=0 t=0

OpenStudy (anonymous):

correct?

OpenStudy (anonymous):

actually am kinda stuck here we have \(t(t-1)(t^2+a)=0 \) and we want to this equation has a one root

OpenStudy (anonymous):

but x=0 is not a root for equations am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you getting it?

OpenStudy (anonymous):

and x=1 is a commen root am i right?

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

see the equation for t

OpenStudy (anonymous):

but no wait

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

what is the answer?

OpenStudy (anonymous):

yes putting 1 in both the eq. a=-1 ans. no. of values of a = 1.

OpenStudy (anonymous):

putting 1 in both the eq. gives a=-2

OpenStudy (anonymous):

1^3+a(1)+1=0 1^4+a(1)^2+1=0 ??

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

see only possible commen roots for 2 equations are t=0 , t=1 or t^2=-a am i right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but t=0 and t^2=-a will not work try to work it out by substituting

OpenStudy (anonymous):

yes